# A numerical study of bounds in the correlations of fractional quantum Hall states

### 28 Feb 2024 v3

@@ I Introduction
In this paper, we investigate this question numerically and provide definitive evidence that weak maximal chirality is insufficient. Thus, we expect only the ideal conformal block wavefunctions to saturate Haldane bound. We compute the static structure factors in the long wavelength limit for FQH states at $\nu=1/3,1/5,\text{and}\ 2/5$ using rotationally invariant two-dimensional Hamiltonians. To this end, we use density matrix renormalization group on an infinite cylinder[15] of circumference $L_{y}$ and approach the 2D-limit by considering large $L_{y}/\ell$. We compute $\bar{S}_{4}$, the coefficient of $\mathcal{O}\left((k\ell)^{4}\right)$ term in the long-wavelength expansion of guiding center static structure factor, and show that it is larger than the Haldane bound for FQH ground states of generic Hamiltonians. We substantiate this observation by analyzing perturbations around the model Hamiltonians of Laughlin FQH states at $\nu=1/3$ and $\nu=1/5$. It is shown that $\bar{S}_{4}$ increases above the bound as the perturbation is introduced.
@@
This paper is organized as follows. In section II, we review some general properties of the guiding center static structure factor that measures only the intra-LL correlations without assuming rotational invariance. A review of the composite-boson formulation, that is used to express the topological quantities, is given in section III. The derivation and statement of the Haldane bound with only translational and inversion symmetries can be found in Ref. [2] and is reproduced in appendix A. It is shown in appendix B that rotational invariance is required for the saturation. Therefore, we specialize to the isotropic FQH states starting in section IV. Our numerical results, showing that neither the Haldane bound is saturated nor the conjectured expression of $\bar{S}_{6}$ is universal for rotationally invariant but weakly maximally chiral fractional quantum Hall ground states, are presented in section V. We substantiate our results by analyzing perturbations around ideal Hamiltonians in section VI. Our findings are summarized alongside remarks on experimental relevance of our results in section VII. Finally, we interpret two of the spectral sum rules derived in Ref. [25] in a purely lowest Landau level formulation in appendix C.
@@ II Guiding center static structure factor
In this section, we review some general properties of the static structure factor of FQH states in the presence of translation and inversion symmetries, without invoking rotational invariance a priori. In appendix B, we show that the saturation of Haldane bound requires rotational invariance, therefore we will specialize to only the isotropic case for numerical investigations. Nevertheless, the following two sections are formulated without this assumption since it is not a requirement for the existence of FQH states or the Haldane bound, and experiments break continuous rotational symmetry. Relaxing this condition has previously led to the discovery of geometric fluctuations as a fundamental degree of freedom in FQH states.[26]
@@
and $C(A,B)$ = $C(B,A)$ = $\langle\tfrac{1}{2}\{A,B\}\rangle$- $\langle A\rangle\langle B\rangle$ is the correlator of two operators (which in general do not commute, although they do in this case), and $C(A^{\dagger},A)$ is real, as in (1), where $S(-\bm{k})$ = $S(\bm{k})$. In a uniform magnetic field, the Faraday tensor in the 2D plane is parameterized by
@@ IV Haldane bound for isotropic FQH states and sixth order term in the static structure factor
$\bar{S}_{4}$ $\geq\frac{\pi\ell^{2}}{\hbar}|\bar{\eta}_{H}|.$ (22)
$\bar{\eta}_{H}$ $=-\frac{\hbar}{8\pi\ell^{2}}\frac{2S_{\rm GC}}{q}$ (23)
$=\frac{\hbar}{8\pi\ell^{2}}\nu(\mathcal{S}-1)$ (24)
where we have assumed that only the lowest Landau level is filled (partially), $eB>0$ and $\nu=p/q$. The guiding center contribution to Hall viscosity $\bar{\eta}_{H}$ is expressed in terms of the guiding center spin of the composite-boson $S_{GC}$ as defined in the section III and comes purely from correlations within the lowest Landau level. Since $S_{GC}$ is odd under particle-hole (PH) symmetry, the expression of Haldane bound is PH symmetric. In the second line, we have expressed it using the conventional Wen-Zee shift $\mathcal{S}=\nu^{-1}N-N_{\phi}$.[32]
@@
A second property of $\bar{S}(\bm{k})$ has been conjectured in the literature using field theoretic approaches.[16, 17, 18, 19, 20] It states that $\bar{S}_{6}$ contains information about the chiral central charge of edge theory and is given by:
$\bar{S}_{6}$ $=\frac{\nu\left(3\bar{\mathcal{S}}^{2}+1\right)-\bar{c}+12\nu\rm{var}(s)}{96}$ (25)
$=\frac{12\nu\left((S_{\rm GC}/p)^{2}+{\rm var}(s)\right)-\bar{c}}{96}$ (26)
where $\bar{c}=c_{-}-\nu$ as defined in Eq. (17) and $c_{-}$ is the chiral central charge of the edge theory defined as the central charge of right-movers minus that of left-movers. It is the quantum anomaly that appears in the Virasoro algebra obeyed by edge state momentum density $\pi_{a}(x)$ as in Eq. (16). The second line is expressed in the composite-boson language. Moreover, ${\rm var}(s)\equiv\frac{s^{T}K^{-1}s}{t^{T}K^{-1}t}-\left(\frac{s^{T}K^{-1}t}{t^{T}K^{-1}t}\right)^{2}$ is an additional quantity conjectured to occur in the case of multi-component Abelian FQH states which has been called the “orbital spin variance”.[35, 36] Here $K$ corresponds to the $K$-matrix of the FQH state with $t$ and $s$ representing the charge and spin vectors respectively[37]. At present, we are not aware of how to interpret the orbital spin variance in the composite-boson language.
We compare this conjecture with our results for the Moore-Read and $\nu=2/5$ Jain-like FQH ground state where we are not aware of any previous numerical calculation for $\bar{S}_{6}$ (although see [23] for an analytical calculation for Moore-Read wavefunction). Importantly, we will show that for generic FQH states, $\bar{S}_{6}$ is smaller than its value in the above equation.
@@ V.3 $\nu=2/5$
In (e), we plot the data for the ground state at $\nu=2/5$ in the presence of $V_{1}$-Haldane pseudopotential interaction. This state is usually interpreted as an integer quantum Hall state of composite-fermions (CFs) at filling fraction $\nu_{\rm cf}=2$,[14] where the CFs are constructed by attaching two flux quanta to electrons. Ref. 43 provided a confirmation of the Jain FQH-like ground state in the infinite cylinder geometry through the computation of its Hall viscosity, chiral central charge, and topological spin. However, we emphasize that the ground state of $V_{1}$ pseudopotential is not identical to Jain wavefunction at $\nu=2/5$, but has the same topological properties.

### 03 Mar 2024 v4

@@ I Introduction
We investigate this question numerically and provide definitive evidence that weak maximal chirality is insufficient. Thus, we expect only the ideal conformal block wavefunctions to saturate Haldane bound. We compute the static structure factors in the long wavelength limit for FQH states at $\nu=1/3,1/5,\text{and}\ 2/5$ using rotationally invariant two-dimensional Hamiltonians. To this end, we use density matrix renormalization group on an infinite cylinder[15] of circumference $L_{y}$ and approach the 2D-limit by considering large $L_{y}/\ell$. We compute $\bar{S}_{4}$, the coefficient of $\mathcal{O}\left((k\ell)^{4}\right)$ term in the long-wavelength expansion of guiding center static structure factor, and show that it is larger than the Haldane bound for FQH ground states of generic Hamiltonians. We substantiate this observation by analyzing perturbations around the model Hamiltonians of Laughlin FQH states at $\nu=1/3$ and $\nu=1/5$. It is shown that $\bar{S}_{4}$ increases above the bound as the perturbation is introduced.
@@
This paper is organized as follows. In section II, we review some general properties of the guiding center static structure factor that measures only the intra-LL correlations without assuming rotational invariance. A review of the composite-boson formulation, that is used to express the topological quantities, is presented in section III. The derivation and statement of the Haldane bound with only translational and inversion symmetries can be found in Ref. [2] and is reproduced in appendix A. It is shown in appendix B that rotational invariance is required for the saturation. Therefore, we specialize to the isotropic FQH states starting in section IV. Our numerical results, showing that neither the Haldane bound is saturated nor the conjectured expression of $\bar{S}_{6}$ is universal for rotationally invariant but weakly maximally chiral fractional quantum Hall ground states, are presented in section V. This conclusion is further substantiated through an analysis of perturbations around ideal Hamiltonians in section VI. Our findings are summarized alongside remarks on experimental relevance in section VII. Lastly, we interpret two of the spectral sum rules derived in Ref. [25] in a purely lowest Landau level formulation in appendix C.
@@ II Guiding center static structure factor
In this section, we review some general properties of the static structure factor of FQH states in the presence of translation and inversion symmetries, without invoking rotational invariance a priori. In appendix B, we show that the saturation of Haldane bound requires rotational invariance, therefore we will specialize to only the isotropic case for numerical investigations. Nevertheless, the following two sections are formulated in a more general setting since it is not a requirement for the existence of FQH states or the Haldane bound, and experiments break continuous rotational symmetry. Moreover, relaxing this condition has previously led to the discovery of geometric fluctuations as a fundamental degree of freedom in FQH states.[26]
@@
where $N$ is the number of particles, $C(A,B)$ = $C(B,A)$ = $\langle\tfrac{1}{2}\{A,B\}\rangle$- $\langle A\rangle\langle B\rangle$ is the correlator of two operators (which in general do not commute, although they do in this case), and $C(A^{\dagger},A)$ is real, as in (1), where $S(-\bm{k})$ = $S(\bm{k})$. In a uniform magnetic field, the Faraday tensor in the 2D plane is parameterized by
@@ IV Haldane bound for isotropic FQH states and sixth order term in the static structure factor
$\bar{S}_{4}$ $\geq\pi\left|\frac{\bar{\eta}_{H}}{eB}\right|.$ (22)
$\bar{\eta}_{H}$ $=-\frac{eB}{4\pi}\frac{S_{\rm GC}}{q}$ (23)
$=\frac{eB}{8\pi}\nu(\mathcal{S}-1)$ (24)
where we have assumed that only the lowest Landau level is filled (partially), and $\nu=p/q$. The guiding center contribution to Hall viscosity $\bar{\eta}_{H}$ is expressed in terms of the guiding center spin of the composite-boson $S_{GC}$ as defined in the section III and comes purely from correlations within the lowest Landau level. Since $S_{GC}$ is odd under particle-hole (PH) symmetry, the expression of Haldane bound is PH symmetric. In the last line, we have expressed it using the conventional Wen-Zee shift $\mathcal{S}=\nu^{-1}N-N_{\phi}$.[32]
@@
A second property of $\bar{S}(\bm{k})$ has been conjectured in the literature using field theoretic approaches.[16, 17, 18, 19, 20] It states that $\bar{S}_{6}$, for chiral FQH states, contains information about the chiral central charge of edge theory and is given by:
$\bar{S}_{6}$ =±$=\pm\frac{\nu\left(3\bar{\mathcal{S}}^{2}+1\right)-\bar{c}+12\nu\rm{var}(s)}{96}$ (25)
=±$=\pm\frac{12\nu\left((S_{\rm GC}/p)^{2}+{\rm var}(s)\right)-\bar{c}}{96}$ (26)
where $\bar{c}\equiv c_{-}-\nu$ is odd under PH transformation and $c_{-}$ is the chiral central charge of the edge theory defined as the central charge of right-movers minus that of left-movers. It is the quantum anomaly that appears in the Virasoro algebra obeyed by edge state momentum density $\pi_{a}(x)$ as in Eq. (16). Moreover, ${\rm var}(s)\equiv\frac{s^{T}K^{-1}s}{t^{T}K^{-1}t}-\left(\frac{s^{T}K^{-1}t}{t^{T}K^{-1}t}\right)^{2}$ is an additional quantity conjectured to occur in the case of multi-component Abelian FQH states which has been called the “orbital spin variance”.[35, 36] Here $K$ corresponds to the $K$-matrix of the FQH state with $t$ and $s$ representing the charge and spin vectors respectively[37]. The combination $\nu\left((S_{\rm GC}/p)^{2}+{\rm var}(s)\right)=(s-t/2)^{T}K^{-1}(s-t/2)$ is odd under PH transformation. At present, we are not aware of its interpretation in the composite-boson framework. The $\pm$ sign is included to make $\bar{S}_{6}$ even under PH transformation. Chiral FQH states and their PH conjugates correspond to $+$ and -- signs respectively.
We compare this conjecture with our results for the Moore-Read and $\nu=2/5$ Jain-like FQH ground state where we are not aware of any previous numerical calculation for $\bar{S}_{6}$ (although see [23] for an analytical calculation for Moore-Read wavefunction). Importantly, we will show that for weak maximally chiral FQH states, $\bar{S}_{6}$ is bounded above by the theoretical prediction.
@@ V.3 $\nu=2/5$
In (e), we plot the data for the ground state at $\nu=2/5$ in the presence of $V_{1}$-Haldane pseudopotential interaction. This state is usually interpreted as an integer quantum Hall state of composite-fermions (CFs) at filling fraction $\nu_{\rm cf}=2$,[14] where the CFs are constructed by attaching two flux quanta to electrons. Ref. 43 provided a confirmation of the Jain FQH-like ground state in the infinite cylinder geometry through the computation of its Hall viscosity, chiral central charge, and topological spin. Although the ground state of $V_{1}$ pseudopotential is not identical to Jain wavefunction at $\nu=2/5$, it has the same topological properties.